# BrainLeaks: On the Privacy-Preserving Properties of Neuromorphic Architectures against Model Inversion Attacks

### v1

@@ 1 Introduction
  • We first propose BrainLeak-v1, as a spike-based model inversion attack that generates spike-compatible inverted input with introducing surrogate gradient into the spiking domain.
@@
To our knowledge, this is the first rigorous investigation of model inversion vulnerabilities in SNNs. The proposed attacks shed light on the privacy risks and advantages of neuromorphic architectures, motivating further research into data privacy for this emerging neural computing paradigm. We make our code open source for reproducibility. [^1] The code is available in the supplementary material
@@ 3.2 BrainLeaks-v1
i) Gradient Quantization: In this first phase, the absolute values of the gradients are normalized to the range [0,1]. Then, we proceed to a linear quantization as follows:
$Q_{i}=\mathbbm{1}(Norm(|\nabla_{X_{i}}|)-\eta)$ (3)
Where $\mathbbm{1}(p)$ is the indicator function which takes 1 if $p\geq\eta$ and 00 elsewhere. $Norm$ identifies as normalization by the maximum value of the vector, and $\eta$ is an empirical quantization threshold.
ii) Sign Extraction: The original gradient signs are combined with the sampled binary mask. This imposes the directionality of the gradients onto the mask, forming ternary discrete gradients of -1,0,1.
iii) Overflow-Aware Transformation: The ternary gradients are used to directly update the input spikes. This update might potentially change spike values to fall outside the valid range of 0, 1, leading to invalid values of -1 or 2. Therefore, we clip the updated spikes, converting out-of-range values (-1 and 2 to 0 and 1, respectively) and keeping valid spike values (0 and 1) unchanged.

### v2

@@ 1 Introduction
  • We first propose BrainLeak-v1, as a spike-based model inversion attack that projects the continuous surrogate gradients obtained during backpropagation into the discrete spiking domain, enabling the generation of spike-compatible inverted inputs.
@@
To our knowledge, this is the first rigorous investigation of model inversion vulnerabilities in SNNs. The proposed attacks shed light on the privacy risks and advantages of neuromorphic architectures, motivating further research into data privacy for this emerging neural computing paradigm. We make our code open source for reproducibility.
@@ 3.2 BrainLeaks-v1
i) Gradients Binarization: In this initial phase, the absolute values of the gradients are normalized to the range $[0,1]$. Subsequently, we derive binerized gradients $G_{i}$ as follows:
$G_{i}=Bernoulli\left(Norm\left|\nabla_{X_{i}}\right|\right)$ (3)
Here, $Bernoulli(p)$ refers to the Bernoulli distribution, which produces binary outputs based on the probability $p$, and $Norm$ identifies as normalization by the maximum value of the vector.
ii) Directional Fusion: The original gradient signs are combined with the sampled binary mask. This incorporates the directionality of the gradients onto the mask, forming ternary discrete gradients of -1,0,1.
iii) Overflow Thresholding: The ternary gradients are used to directly update the input spikes. This update might potentially change spike values to fall outside the valid set of ${0,1}$, leading to invalid values of -1 or 2. Therefore, we clip the updated spikes, converting out-of-bounds values (-1 and 2 to 0 and 1, respectively) and keeping valid spike values (0 and 1) unchanged.